### v1

# The nature of mathematical models

Abstract
Modeling has become a widespread, useful tool in mathematics applied to diverse fields, from physics to economics to biomedicine. Practitioners of modeling may use algebraic or differential equations, to the elements of which they attribute an intuitive relationship with some relevant aspect of reality they wish to represent. More sophisticated expressions may include stochasticity, either as observation error or system noise. However, a clear, unambiguous mathematical definition of what a model is and of what is the relationship between the model and the real-life phenomena it purports to represent has so far not been formulated. The present work aims to fill this gap, motivating the definition of a mathematical model as an operator on a Hilbert space of random variables, identifying the experimental realization as the map between the theoretical space of model construction and the computational space of statistical model identification, and tracing the relationship of the geometry of the model manifold in the abstract setting with the corresponding geometry of the prediction surfaces in statistical estimation.
1 Introduction
Mathematical modeling is central to defining and solving problems in the physical sciences, engineering or economics. Over the past few decades it has found progressively more widespread applications also in biomedicine and other “soft” sciences. In the following we will make use, just by way of example, of naively simple problems in biology: while this choice stems from the background of the author as a biomedical modeler, the mathematical ideas are equally applicable to any of the above application fields.
When considering a concrete problem in any field, the mathematical modeler is usually confronted with two distinct tasks: the first consists of imagining a potentially useful functional relationship among quantities, amenable to analysis and comparison with the real-world (the modeling phase proper); the second consists in attempting to quantify the model parameters, leading to some hopefully good representation of the main features of the considered real-world phenomena (the model identification phase). In practice, these tasks may succeed each other cyclically, an abstract formulation leading to an identification attempt which, if reputed unsuccessful, leads to a further formulation and so on.
The initial, modeling phase thus consists in imagining possible relationships among some quantitative features of interest of the reality under study, relationships usually expressed as algebraic or differential equations of various kinds. The initial concern of the present work is to understand exactly what the variables in these equations represent. Suppose for example we wished to estimate weight from height, and we were to write, as a very first attempt to quantify the relationship of weight as a function of height, the equation
$W=kH$ (1)
Notice that at the present stage we are not dealing with any numbers at all, we are at the first stage of idealization of a relationship between a generic concept of height with a generic concept of weight. Our basic question is then: what, in the above equation, is $H$? The immediate, naive answer would be "Height!". Yes, clearly, but whose height? Obviously $H$ is not meant to represent the numerical value of the height of any single individual, to be put in relationship with the weight $W$ of that same individual: changing individuals would then in general change the value of the parameter $k$, which is instead intended to express some general proportion between the height of people and their weight. Similarly, Eq. 1 does not express the relationship between an algebraic variable $W$ and an algebraic variable $H$: we do not imagine people’s weights to follow exactly that functional relationship. Rather, we might consider $H$ as indicating a quantity (height) which could be measured on some individual or on a group of individuals, quite possibly with variability due to heterogeneity among people and/or to measurement error: while we have not yet done so at the present stage of the modeling exercise, we could do so in the future. We are thus led to considering $H$ as a random variable.
At this point we consider what $W$ might be. Notice that, if we were to consider hypothetically the observable weight of individuals, and, following the previous line of reasoning, we were to consider such observable weight of individuals as a random variable, this random variable (let us denote it provisionally by $W^{o}$) would not be the same as $W$. That $W$ is a random variable follows from the fact that it is a function of some numerical value $k$ and of another random variable $H$, so yes, $W$ is indeed a random variable. But $W$ is that random variable obtained from $H$ through some sort of mathematical manipulation, whereas $W^{o}$ is another random variable, representing some characteristic of the considered people, which may or may not depend on $H$ and even if it did it may well in general not depend on it through Eq. 1. In other words, we might consider $W^{o}$ as the objective (hence the o) random variable that we are trying to approximate through some other random variable $W$, which is itself computable from $H$. We are thus led to the definition of a model, which will be formalized in the following, as an operator yielding a random variable (in this case $W$) with which we try to approximate another, objective random variable(in this case $W^{o}$). As will be made more precise in the following, approximation has meaning within a space endowed at least with a norm, while other useful concepts like projections will need further structure (inner products), thus leading us to deal with of a Hilbert space of random variables.
If we now consider the second, identification phase of modeling practice, we see that it is commonly performed through the minimization of a cost functional (e.g. by Ordinary Least Squares, OLS). Identification of course depends on having sampled some individuals from the population of interest and having measured on them both the intended predictor variable(s) and the intended criterion or target variable. This procedure, which is inherently computational, is executed within an $n$-dimensional Euclidean space, referred to as the case space.
While both phases of the modeling process, model building and model identification, have been and are conducted innumerable times by modelers over centuries and all over the world, the relationship between ideal model building and concrete statistical estimation computations has been a matter the "obvious" intuition of the correspondence between the two. The aim of the present work is to make formally precise the relationship between the ideal construction of a mathematical model in the space of random variables and the statistical estimation procedure in $n$-dimensional case space. It should be noted that the consideration of a Hilbert space of random variables is not new, see for example Small [12]. However, the subsequent development has centered on the characterization of conditional expectation as the projection onto a linear subspace [4]. The present work, instead, focuses on the modeling process itself, giving rise to a generally nonlinear model manifold, and on the relationship of this construction with the (geometrical interpretation of) commonly performed statistical parameter estimation procedures. There is however a closer link with the whole field of information geometry [9, 11, 6], dealing with manifolds of probability distributions, in that there exists a correspondence between (equivalence classes of) random variables and their induced probability distributions.
In the following Methods section notation is first established. Then, supporting (well known) theorems are recalled and the formal constructions of a model, a model function and a model manifold are introduced. The geometry of statistical estimation, in the simple OLS case, is recalled. The problem is then tackled of how the challenges introduced by model nonlinearity in the two settings make the introduction of appropriate tangent spaces necessary. The mapping of the geometry in the random variable space to the geometry of the statistical estimation case space is finally formally defined. A brief Discussion and Conclusions section recaps the results obtained, clarifies the relationship between model manifolds, conditional expectation and information geometry and points to possible further developments.
2 Methods
2.1 Notation
2.1.1 Standard notation
Scalars are represented as (Latin or Greek) lower-scale plain-face letters ($a$, $\alpha$), vectors as lower-scale bold letters ($\boldsymbol{b}$, $\boldsymbol{\beta}$). However, when treating generic quantities, which may be either vector- or scalar-valued, we will often use plain-face rather than bold characters.
Matrices and higher-order arrays are written as (capital) boldface latin letters ($\boldsymbol{H}$ or $H$ when clear from the context). An array may show its dimensions as subscripted integers separated by a cross ($A_{n\times m}$ is the matrix $A$ with $n$ rows and $m$ columns); an array element is identified by its indices (row $r$, column $c$) as $a_{r,c}$ or $a_{rc}$ when no confusion is possible. The $r^{th}$ row and $c^{th}$ column of the matrix $A$ are indicated respectively as $a_{r.}$ and $a_{.c}$   . Notice the difference between $\boldsymbol{x}_{i}$ (the $i^{th}$ vector of a sequence), and $x_{i}$ (the $i^{th}$ scalar of a sequence, usually the $i^{th}$ element of vector $\boldsymbol{x}$).
In order to distinguish between a random variable and its realization, we typically indicate the (scalar- or vector- valued) random variable upper case letter ($U$, $\boldsymbol{U}$) and its realization as the appropriate lower case letter ($u$, $\boldsymbol{u}$).
In order to keep notation consistent we will strive to have indices and dimensions correspond throughout, with $i=1,\dots,n$ referring to cases (observations, criterion values, dependent variable values), $j=1,\dots,q$ referring to parameters, and $k=1,\dots,m$ referring to predictors (independent variables).
The real numbers are denoted by $\mathbb{R}$. Positive reals are denoted by $\mathbb{R}^{+}$. Spaces or manifolds are generally indicated by script capital letters like $\mathcal{V}$, $\mathcal{H}$ or $\mathcal{M}$.
Transposition of an array is denoted by superscript capital $T$, like in $J^{\intercal}$. An array, of appropriate dimensions, having all elements equal to 00 (or equal to $1$) is indicated as $\boldsymbol{0}$ (or as $\boldsymbol{1}$ ). The identity matrix of appropriate dimensions is indicated by $\boldsymbol{I}$
The space spanned or generated by a set of vectors is indicated as $\mathcal{S}({\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{m}})$ ; when applied to a matrix, the span function identifies the space generated by the column vectors of the matrix, $\mathcal{S}(A)$
The null space or kernel of a matrix $\boldsymbol{A}$ is denoted by $Ker(\boldsymbol{A})$: $Ker(\boldsymbol{A})=\{\boldsymbol{v}:\boldsymbol{Av}=0\}$.
The orthogonal complement to a linear subspace $V$ is denoted by $V^{\perp}$ ("V-perp"): $V^{\perp}=\{\boldsymbol{x}:\boldsymbol{x}^{\intercal}\cdot\boldsymbol{v}=0\ \ \forall\boldsymbol{v}\in V\}$.
The generalized inverse of a matrix $\boldsymbol{A}$ is indicated as $\boldsymbol{A}^{-}$ and satisfies $\boldsymbol{A}\boldsymbol{A}^{-}\boldsymbol{A}=\boldsymbol{A}$
Probability measure is indicated by a capital $P$, while the probability density function (pdf) of a random variable is indicated by a lower case $p$ ; in this way, the probability of an event $A$ is
[…]
and if $B$ is a measurable set in $\mathbb{R}$ with respect to the usual Borel $\sigma$-algebra $\mathcal{B}$, and $X$ is a random variable from $\Omega$ to $\mathbb{R}$ with density $p_{X}(x)$ then
[…]
Under these conditions the expectation of a function $g(x)$ of the random variable $X$ is
[…]
Often used abbreviations are   WRT   (With Respect To) and   WLOG   (Without Loss of Generality).
End of proofs are indicated with the square box ∎.
2.1.2 Further notation
So far, the notation has been following standard statistical practice. In what follows, however, some necessary further notational conventions are established, which will make it easier to develop the argument in a precise way.
We let $\boldsymbol{x}^{o}=[x^{o}_{i}]_{n\times 1}$ be a vector of observed ("o" stands for observed) realizations of a dependent variable (criterion), possibly corresponding to arbitrarily assigned or anyhow measured values of $m$ independent variables (predictors) $\boldsymbol{U}=[u_{ik}]_{n\times m}=[\boldsymbol{u}_{i.}]_{n\times m}$ , one of which is often time $t$. We also let $\boldsymbol{\theta}=[\theta_{j}]_{q\times 1}$ be a (vector-valued) parameter, $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{q}$, and we suppose that the (generally nonlinear) functional form of the relationship between criterion and predictors is
$x^{o}_{i}=x(\boldsymbol{u}_{i},\boldsymbol{\theta})+\varepsilon_{i}$
or
$\boldsymbol{x}^{o}=\boldsymbol{x}(\boldsymbol{U},\boldsymbol{\theta})+\boldsymbol{\varepsilon}$
where $\boldsymbol{\varepsilon}$ is an $n\times 1$ vector of "errors", i.e. of casually, randomly occurring differences between modeled and observed values of the criterion for the given predictors. Notice that while $\boldsymbol{x}^{o}$ is a column array of values, $x()$ is a scalar function and $\boldsymbol{x}()$ is a vector-valued function.
In much of our discussion we will not need to indicate the predictors explicitly, since they are part of the experimental design and are fixed. We will thus simply write our predictor function as
$\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta})\quad\left(=\boldsymbol{x}(\boldsymbol{U},\boldsymbol{\theta})\right)$
We indicate with a "hat" the estimated value of the parameter, $\hat{\boldsymbol{\theta}}$, to which corresponds a "best" prediction or forecast $\hat{\boldsymbol{x}}=\boldsymbol{x(\hat{\boldsymbol{\theta}})}$. We may also suppose that, the postulated relationship being the true one, there exists a "true" value of the parameter, $\boldsymbol{\theta}^{*}$, to which there corresponds the true state of the system $\boldsymbol{x}^{*}$ if it were observed without error.
For any value of the parameter we may compute the first derivative of the forecast with respect to the parameter, the Jacobian:
$\boldsymbol{J}=\left[\frac{\partial x_{i}(\boldsymbol{\theta})}{\partial\theta_{j}}\right]_{n\times q}$
and, more specifically, we compute the Jacobian at the estimated parameter value as
$\hat{\boldsymbol{J}}=\left[\frac{\partial x_{i}(\hat{\boldsymbol{\theta}})}{\partial\theta_{j}}\right]_{n\times q}=\left.\frac{\partial x_{i}(\boldsymbol{\theta})}{\partial\theta_{j}}\right|_{\hat{\boldsymbol{\theta}}}$
Using an alternative notation for the Jacobian we may write (in the spirit of Seber [2])
$\boldsymbol{\overset{.}{x}}=\boldsymbol{J}=\frac{\partial x_{i}(\boldsymbol{\theta})}{\partial\theta_{j}}$
so that the three-dimensional array of second-order partial derivatives is denoted as
$\boldsymbol{\overset{..}{x}}=\left[\frac{\partial^{2}x_{i}(\boldsymbol{\theta})}{\partial\theta_{j}\ \partial\theta_{k}}\right]_{q\times n\times q}$
with the obvious meaning being attributed to $\hat{\boldsymbol{\overset{.}{x}}}$, $\hat{\boldsymbol{\overset{..}{x}}}$, $\boldsymbol{\overset{.}{x}}^{*}$ and $\boldsymbol{\overset{..}{x}}^{*}$
Multiplications involving these three-dimensional arrays are defined, for the scope of these notes, simply as the slice-wise equivalent: imagine the array $\boldsymbol{A}_{m\times n\times q}$ to be a pile of $n$ slices, each of which is a $m\times q$ matrix $(\boldsymbol{A}_{.i.})_{m\times q}$ Then pre-multiplication by a matrix $\boldsymbol{B}_{r\times m}$ and post-multiplication by a matrix $\boldsymbol{C}_{q\times s}$ are defined as producing $n$-piles of matrices $(\boldsymbol{BA}_{.i.})_{r\times n\times q}$ or $(\boldsymbol{A}_{.i.}\boldsymbol{C})_{m\times n\times s}$ respectively, with the obvious understanding that dimensions collapsing to $1$ disappear, so that, for example, we treat $\boldsymbol{v}_{1\times m}\boldsymbol{A}_{m\times n\times q}$ simply as the matrix $(\boldsymbol{vA})_{n\times q}$.
  • a bold symbol indicates a vector. Thus $\boldsymbol{\varepsilon}$, $\boldsymbol{\xi}$ and $\boldsymbol{e}$ are vectors.
  • an orthonormal basis for $\mathbb{R}^{n}$ is a collection of $n$ unit-length, orthogonal basis vectors, say $\boldsymbol{H}=\{\boldsymbol{\eta}_{1},\dots,\boldsymbol{\eta}_{n}\}$ or $\boldsymbol{P}=\{\boldsymbol{\rho}_{1},\dots,\boldsymbol{\rho}_{n}\}$.
  • a vector $\boldsymbol{v}$ is the sum of its components $\{\boldsymbol{v}_{i}\}_{i=1}^{n}$, each of which can be expressed as a coefficient multiplying a basis vector, say $\boldsymbol{v}_{i}=v_{\eta}^{i}\cdot\boldsymbol{\eta}_{i}$, where we emphasize that the coefficient $v_{\eta}^{i}$ refers to the chosen basis $\boldsymbol{H}$
    (Notice that we are NOT using here Einstein’s summation notation, the dot above denotes scalar multiplication in the vector space $\mathbb{R}^{n}$ over the reals)
  • we may collect the coefficients into a column array of coefficients, relative to the chosen basis, say $\boldsymbol{v}_{\eta}=\{v^{1}_{\eta},...,v^{n}_{\eta}\}^{T}$    or    $\boldsymbol{v}_{\rho}=\{v^{1}_{\rho},...,v^{n}_{\rho}\}^{T}$
  • we can thus concisely express a vector as its expansion with respect to a given basis:    $\boldsymbol{v}=\boldsymbol{E}\boldsymbol{v}_{\eta}=v_{\eta}^{i}\boldsymbol{\eta}_{i}$    or    $\boldsymbol{v}=\boldsymbol{P}\boldsymbol{v}_{\rho}=v_{\rho}^{i}\boldsymbol{\rho}_{i}$
    (Notice that here we ARE using Einstein’s summation notation)
2.2 Geometry of Estimation in Case Space ($\mathbb{R}^{n}$)
We summarize here the standard geometrical interpretation of Ordinary Least Squares parameter estimation. While none of the material in this subsection is new (e.g. see Seber [1, 2]), the following description should help to clarify the correspondence of this with subsequently presented material.
When talking about modeling in an applied setting we are typically discussing the “usual representation”
$x=x(u,\boldsymbol{\theta})$ (2)
where a dependent variable $x$ is shown to depend upon some independent predictor $u$ (often indicating time in dynamical systems) via some functional form and some $q$-dimensional parameter $\boldsymbol{\theta}$. This is typically graphed showing curves for $x$ as a function of $u$, different curves corresponding to different values of $\boldsymbol{\theta}$. Since by changing the parameter value the prediction $x(u,\boldsymbol{\theta})$ obviously changes, this clearly leads to thinking about estimation as adapting the predicted curve by minimizing some functional of the "errors" (differences between predicted and observed values of $x$), yielding    $\hat{\boldsymbol{\theta}}$    as the $argmin$ of the functional.
However the values $u$ takes are fixed (by design or by sampling), and we may drop $u$ from notation, since what we are really interested in is the relationship between parameter and prediction. Also, we may indicate with $\boldsymbol{x}$ the set of predicted values that $x$ takes for each of $n$ experimental units (corresponding to a given set of predictors) as $\boldsymbol{\theta}$ varies:
$\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta})$ (3)
and we may also indicate with $\boldsymbol{x}^{o}$ (observed) the actually measured or observed values of $x$ in the $n$ experimental units. We find thus useful to work in ($n$-dimensional) $\boldsymbol{x}$-space or case space, $\mathbb{R}^{n}$. Case space is therefore defined as the $n$-dimensional Euclidean space where each axis represents one of the experimental units, and where along each axis we may identify both the actually observed value for that unit and any predicted value for that unit (depending on the chosen $\boldsymbol{\theta}$ parameter value). In other words, case space is the $n$-dimensional space where all possible values of $\boldsymbol{x}$, either measured or predicted, referring to the $n$ xperimental units can be represented. The appeal of working in case space consists in an immediate visualization of the geometry of the problem: the whole observed sample realization is a single point $\boldsymbol{x}^{o}$ in $n$-dimensional case space, and we may trace a (generally nonlinear) $q$-dimensional prediction surface $\mathcal{S}$ as the set of values that the predicted $\boldsymbol{x}$ can take as $\boldsymbol{\theta}$ varies. Indeed, if $\boldsymbol{\theta}\in\mathbb{R}^{q}$ and $x(\boldsymbol{\theta})$ is sufficiently regular then the prediction surface $\mathcal{S}$ is a $q$-dimensional manifold embedded in case space ($\mathbb{R}^{n}$), having as a global chart map the canonical map $\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta})\in\mathbb{R}^{n}\rightarrow\boldsymbol{\theta}\in\mathbb{R}^{q}$, everywhere.
A number of useful remarks may be offered at this point:
  • $\mathbb{R}^{n}$ has the structure of a vector space over the reals
  • we can define on $\mathbb{R}^{n}$ the Standard Euclidean topology $\mathcal{O}$
  • $\mathcal{O}$ generates the Borel $\sigma$-algebra $\mathcal{B}$
  • $(\mathbb{R}^{n},\mathcal{B})$ is a measurable space
  • we have the Lebesgue measure $\lambda$ defined on $\mathcal{B}$ with values in +:=$\mathbb{R}^{+}:=\{r\in\mathbb{R},r\geqslant 0\}$, making $(\mathbb{R}^{n},\mathcal{B},\lambda)$ into a measure space
  • we can endow $\mathbb{R}^{n}$ with a norm, the usual euclidean norm $|\boldsymbol{x}|:=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$
  • with respect to $|\boldsymbol{x}|$ the vector space $\mathbb{R}^{n}$ is complete, hence Banach
  • we can define on $\mathbb{R}^{n}$ an inner product compatible with the norm, $\langle\boldsymbol{x},\boldsymbol{y}\rangle:=\sqrt{\sum_{i=1}^{n}x_{i}y_{i}}$, making $\mathbb{R}^{n}$ Hilbert
2.2.1 Linear relationship in case space: geometry
We begin by hypothesizing that the relationship given by Eq.3 between $\boldsymbol{x}$ and $\boldsymbol{\theta}$ is actually linear. Referring to Fig.1 at page 1 we can now identify a number of relevant geometric features of the problem:
  • ${\boldsymbol{\theta}}$ is a $q$-dimensional parameter
  • $\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta})$ is a function mapping (in this case linearly) the set   $\Theta$   of allowable parameter values onto $\mathbb{R}^{n}$, $\boldsymbol{x}:\Theta\rightarrow\mathbb{R}^{n}$
  • $\mathcal{S}$ (the prediction surface) is the set of possible values taken by the function $\boldsymbol{x}(\boldsymbol{\theta})$; in this linear case it is a $q$-dimensional subspace of $\mathbb{R}^{n}$
  • since we assume by hypothesis the relationship given by Eq.3, then $\boldsymbol{x}^{*}$ is the state, the supposedly true value of the observed phenomenon, necessarily lying on the supposedly true expectation surface, at a position indexed by the supposedly true parameter value $\boldsymbol{\theta}^{*}$
  • $\boldsymbol{x}^{o}$ is the observation, the observed point in case space, assumed to be generated by the true linear relationship, but to be affected by some observation error
  • $\hat{\boldsymbol{x}}$ is our best estimate for the state, the point on $\mathcal{S}$ closest to the observation $\boldsymbol{x}^{o}$, indexed by the parameter value $\hat{\boldsymbol{\theta}}$; notice that it makes sense of speaking about closeness since, as mentioned before, $\mathbb{R}^{n}$ is normed; also, we are arbitrarily taking minimum distance between $\boldsymbol{x}^{o}$ and $\mathcal{S}$ as the basis for our estimation (OLS)
  • $\boldsymbol{{\varepsilon}}=\boldsymbol{x}^{o}-\boldsymbol{x}^{*}$ is the error, the difference between the observation and the true state; since we do not know $\boldsymbol{x}^{*}$ we do not know $\boldsymbol{\varepsilon}$, even though we might make hypotheses about how it should be distributed
  • $\boldsymbol{e}=\boldsymbol{x}^{o}-\hat{\boldsymbol{x}}$ is the residual, the difference between the observation $\boldsymbol{x}^{o}$ and our best guess $\hat{\boldsymbol{x}}$ given the observation and the assumed relationship
  • ${\boldsymbol{\xi}}=\hat{\boldsymbol{x}}-\boldsymbol{x}^{*}$ is the flaw (newly introduced terminology, a name for this object could not be found in the standard literature), the assessment mistake we would make by placing our best guess at $\hat{\boldsymbol{x}}$ instead of at $\boldsymbol{x}^{*}$
Figure 1: The geometry of a linear relationship in case space
2.2.2 Linear relationship in case space: estimation by projection
Given the above geometry, we can now describe a possible procedure for estimating $\boldsymbol{\theta}$. Notice first that, since $\mathcal{S}$ is a (nonempty) subspace of $\mathbb{R}^{n}$, it is convex and a Hilbert space in its own right. From the Hilbert projection theorem[13] a unique projection operator $\mathcal{P}_{\mathcal{S}}:\mathbb{R}^{n}\rightarrow\mathcal{S}$ can be defined. In fact, if it were $\boldsymbol{x}=\boldsymbol{\theta}U$ for some matrix $U$, we could write $\mathcal{P}_{\mathcal{S}}=U(U^{T}U)^{-1}U^{T}$. Since this projection operator satisfies (always by Hilbert projection theorem)
$\|\boldsymbol{x}-\mathcal{P}_{\mathcal{S}}\boldsymbol{x}\|<\|\boldsymbol{x}-\boldsymbol{y}\|\,,\,\forall\boldsymbol{y}\in\mathcal{S},\boldsymbol{y}\neq\mathcal{P}_{\mathcal{S}}\boldsymbol{x}\,,$
we have that $\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}\in\Theta}{\arg\min}(||\boldsymbol{x}(\boldsymbol{\theta})-\boldsymbol{x}^{o}||^{2})$ ,  where $\boldsymbol{\Theta}$ is the admissible set of values of $\boldsymbol{\theta}$.
The projection $\hat{\boldsymbol{x}}=\mathcal{P}_{\mathcal{S}}\boldsymbol{x}^{o}$ of the observed sample $\boldsymbol{x}^{o}$ onto the prediction surface $\mathcal{S}$ thus identifies the unique point $\hat{\boldsymbol{x}}\in\mathcal{S}$ which is closest to $\boldsymbol{x}^{o}$ in the OLS sense. The value $\hat{\boldsymbol{\theta}}$ such that $\hat{\boldsymbol{x}}=\boldsymbol{x}(\hat{\boldsymbol{\theta}})$ is the desired (point) parameter estimate.
Under reasonable assumptions:
  • the relationship is true
  • observations are determined by the true relationship, design variable(s) and the supposedly true parameter value $\theta^{*}$, plus error
  • errors on the observations of the $n$ experimental units are i.i.d. (independent identically distributed), with a normal distribution of zero mean and variance $\sigma^{2}$
we have:
  1. 1.
    the OLS point estimate coincides with the Maximum Likelihood Estimate (with the attending nice properties of efficiency and consistency of the MLE);
  2. 2.
    asymptotic confidence regions of $\hat{\boldsymbol{x}}=\mathcal{P}_{\mathcal{S}}\boldsymbol{x}^{o}$ on $\mathcal{S}$, hence of $\hat{\boldsymbol{\theta}}$ can be obtained [1].
In fact under these hypotheses:
  • $\|\boldsymbol{\varepsilon}\|^{2}=\sum_{i=1}^{n}\,\varepsilon_{i}^{2}$, where $\forall\,i\,\,\,\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ (Pythagoras theorem);
  • under a rotation of the basis for $\mathbb{R}^{n}$ we can obtain a new basis with the first $q$ elements spanning $\mathcal{S}$ and with the remaining $(n-q)$ elements orthogonal to $\mathcal{S}$;
  • the components of $\boldsymbol{\varepsilon}$ in the new basis are again i.i.d $\mathcal{N}(0,\sigma^{2})$ , which follows from the fact that an orthonormal linear transformation of a standard multivariate normal is a standard multivariate normal;
  • since $\boldsymbol{\varepsilon}=\boldsymbol{\xi}+\boldsymbol{e}$, with $\boldsymbol{\xi}\perp\boldsymbol{e}$, we have again by Pythagoras theorem that $\|\boldsymbol{\varepsilon}\|^{2}=\|\boldsymbol{\xi}\|^{2}+\|\boldsymbol{e}\|^{2}=$, where $\|\boldsymbol{\xi}\|^{2}=\sum_{j=1}^{n}\,\xi_{j}^{2}$ and $\|\boldsymbol{e}\|^{2}=\sum_{i=1}^{n}\,e_{i}^{2}$;
  • notice however that with respect to the new basis
    $\xi_{j}=0\,\,\forall j\in\{q+1,\dots,n\}$   and   $e_{i}=0\,\,\forall i\in\{1,\dots,q\}$, while
    $\xi_{j}\sim\mathcal{N}(0,\sigma^{2}),\,j\in\{1,\dots,q\}$   and    $e_{i}\sim\mathcal{N}(0,\sigma^{2}),\,i\in\{q+1,\dots,n\}$
From the above follow important statistical consequences:
  • $\frac{\|\boldsymbol{\varepsilon}\|^{2}}{\sigma^{2}}=\sum_{i=1}^{n}\,\frac{\varepsilon_{i}^{2}}{\sigma^{2}}=\sum_{i=1}^{n}\,\left(\frac{\varepsilon_{i}}{\sigma}\right)^{2}$,   where    $\frac{\varepsilon_{i}}{\sigma}\sim\mathcal{N}(0,1)$,   so
    $\frac{\|\boldsymbol{\varepsilon}\|^{2}}{\sigma^{2}}\sim\chi_{n}$  (Chi-square with $n$ degrees of freedom)
  • $\frac{\|\boldsymbol{\xi}\|^{2}}{\sigma^{2}}=\sum_{j=1}^{n}\,\frac{\xi_{j}^{2}}{\sigma^{2}}=\sum_{j=1}^{q}\,\left(\frac{\xi_{j}}{\sigma}\right)^{2}$,   where   $\frac{\xi_{j}}{\sigma}\sim\mathcal{N}(0,1)$,   so
    $\frac{\|\boldsymbol{\xi}\|^{2}}{\sigma^{2}}\sim\chi_{q}$  (Chi-square with $q$ degrees of freedom)
  • $\frac{\|\boldsymbol{e}\|^{2}}{\sigma^{2}}=\sum_{i=1}^{n}\,\frac{e_{i}^{2}}{\sigma^{2}}=\sum_{i=q+1}^{n}\,\left(\frac{e_{i}}{\sigma}\right)^{2}$,   where   $\frac{e_{i}}{\sigma}\sim\mathcal{N}(0,1)$,   so
    $\frac{\|\boldsymbol{e}\|^{2}}{\sigma^{2}}\sim\chi_{n-q}$  (Chi-square with $(n-q)$ degrees of freedom)
  • $\frac{\|\boldsymbol{\xi}\|^{2}}{\sigma^{2}}/\frac{\|\boldsymbol{e}\|^{2}}{\sigma^{2}}=\frac{\|\boldsymbol{\xi}\|^{2}}{\|\boldsymbol{e}\|^{2}}\sim\textit{F}_{q,(n-q)}$  (Fisher with $q$ and $(n-q)$ degrees of freedom)
and we can compute a $(1-\alpha)$ confidence region for $\|\boldsymbol{\xi}\|$ !
In other words: under the stated hypotheses ($\varepsilon_{i}$ i.i.d. $\sim\mathcal{N}(0,\sigma^{2})$), taking into account the dimension $n$ of case space as well as the dimension $q$ of the prediction surface $\mathcal{S}$, by observing the size of the residual from $\mathcal{S}$, $\|\boldsymbol{e}\|$, we have information on the likely size of the flaw along $\mathcal{S}$, $\|\boldsymbol{\xi}\|$ !!!
2.2.3 Nonlinear relationship in case space
Figure 2: The geometry of a nonlinear relationship in case space
By referring to Fig.2 we see that a construction, similar to what we have provided for the linear case, does not actually hold any more. The fact is that in this context $\mathcal{S}$ is not a subspace of $\mathbb{R}^{n}$ and in general it may not even be convex, hence Hilbert projection theorem does not apply and a unique projection operator $\mathcal{P}_{\mathcal{S}}:\mathbb{R}^{n}\rightarrow\mathcal{S}$ cannot be defined.
We may however look for a value $\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}\in\boldsymbol{\Theta}}{\arg\min}(||\boldsymbol{x}(\boldsymbol{\theta})-\boldsymbol{x}^{o}||)$ ,  where again $\boldsymbol{\Theta}$ is the admissible set of values of $\boldsymbol{\theta}$, same as for the linear case.
The value $\hat{\boldsymbol{\theta}}$ such that $\hat{\boldsymbol{x}}=\boldsymbol{x}(\hat{\boldsymbol{\theta}})$, the desired (point) parameter estimate, is typically found numerically, by minimization of $\|\boldsymbol{e}(\boldsymbol{\theta})\|$.
We would like to be able to make at least approximate confidence statements about the parameter estimate in the nonlinear case as well. To do so, and in the case that the (local) minimum we have found is isolated, we might wish to use a local linear approximation $T_{\hat{\boldsymbol{x}}}\mathcal{S}$ to the prediction surface around the estimated state $\hat{\boldsymbol{x}}$. Notice that $T_{\hat{\boldsymbol{x}}}\mathcal{S}$ is not in general a subspace, but rather an affine space.
The Hilbert projection theorem however extends naturally to affine subspaces, the key observation being that the problem can be reduced to projecting onto a linear subspace by translating the space appropriately. The projection onto an affine subspace is unique and satisfies an orthogonality condition relative to the associated direction subspace, see Appendix 5.1 for details. In this way, if an isolated minimum of $\|\boldsymbol{e(\boldsymbol{\theta})}\|$ at $\hat{\boldsymbol{x}}$ exists, then we can define a (locally unique) projection operator    $\mathcal{P}_{T_{\hat{\boldsymbol{x}}}\mathcal{S}}:\mathbb{R}^{n}\rightarrow T_{\hat{\boldsymbol{x}}}\mathcal{S}$,    $\mathcal{P}_{T_{\hat{\boldsymbol{x}}}\mathcal{S}}:\boldsymbol{x}^{o}\mapsto\hat{\boldsymbol{x}}$
Under the same assumptions that we made for the linear case:
  • the relationship is true
  • observations are determined by the true relationship, design variable(s) and the supposedly true parameter value $\theta^{*}$, plus error
  • errors on the observations of the $n$ experimental units are i.i.d. (independent identically distributed), with a normal distribution of zero mean and variance $\sigma^{2}$
we have that [2, 3]:
  1. 1.
    there is no guarantee that the local OLS point estimate is the global optimum.
  2. 2.
    approximate asymptotic confidence regions of $\hat{\boldsymbol{x}}$ on $T_{\hat{\boldsymbol{x}}}\mathcal{S}$, hence of $\hat{\boldsymbol{\theta}}$ can be obtained: such regions would not be too different from the correct asymptotic confidence regions of $\hat{\boldsymbol{x}}$ on $\mathcal{S}$ if the curvature of $\mathcal{S}$ at $\hat{\boldsymbol{x}}$ is not too large [14].
3 Geometry of Estimation in the Hilbert space $\mathcal{H}$ of random variables
3.1 Preliminaries
From what has been presented in the Introduction, modeling appears to be the act of hypothesizing mathematical relationships among idealized quantities, some of which may be observed (i.e. numerically measured in a certain number of cases). To make this idea precise, a few relevant, very standard results are listed in the following. Proofs can be found in any textbook of functional analysis or mathematical statistics [15, 16].
First of all, the notion of a quantity that might be observed in a certain number of cases (with possibly different outcomes) coincides with the mathematical concept of a random variable.
Random variables
Let $(\Omega,\mathcal{F},P)$ be a probability space.
Recall that a $n$-dimensional random variable is a measurable map from the probability space $(\Omega,\mathcal{F},P)$ to the measurable space $(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}))$, where $\mathcal{B}(\mathbb{R}^{n})$ is the Borel $\sigma$-algebra of $\mathbb{R}^{n}$.
The Doob-Dynkin lemma for random variables (also see Fig.3)
Theorem (Doob-Dynkin).
Let $(\Omega,\mathcal{F},P)$ be a probability space, $(\mathbb{R}^{n},\mathcal{B}_{n})$ and $(\mathbb{R}^{m},\mathcal{B}_{m})$ measurable spaces, $U:\Omega\rightarrow\mathbb{R}^{n}$ a random variable. Then, a function $X:\Omega\rightarrow\mathbb{R}^{m}$ is $\sigma(U)$-$\mathcal{B}_{m}$-measurable iff $\,\,\exists\,x:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, $g$ $\mathcal{B}_{n}$-$\mathcal{B}_{m}$-measurable, such that $X=x\circ U=x(U)$. Consequently, $X$ is also a random variable.
[…]
Figure 3: By the Doob-Dynkin lemma this diagram commutes
The space $\mathcal{L}^{p}(\Omega,\mathcal{F},\mathbb{P})$
[…]
We define the equivalence class of a measurable function $X$ under the equivalence relationship =$\overset{a.e.}{=}$ as $[X]\in\mathcal{L}^{p}/\overset{a.e.}{=}$ as $[X]=\{Y\in\mathcal{L}^{p}:Y\overset{a.e.}{=}X\}$.
[…]
By an abuse of notation we often call $\mathcal{L}^{p}$ the quotient space $\mathcal{L}^{p}/\overset{a.e.}{=}$ and say that $X\in\mathcal{L}^{p}$ when we really mean $[X]\in\mathcal{L}^{p}/\overset{a.e.}{=}$.
It can finally be shown that $\mathcal{L}^{p}(\Omega,\mathcal{F},P))$, $p\geq 1$ is a Banach space (i.e complete and normed).
The Hilbert space of finite-variance random variables, $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathcal{F},\mathbb{P})$
[…]
Notice that $\mathcal{L}^{p}(\Omega,\mathcal{F},\mathbb{P})$ is not a Hilbert space for $p\neq 2$: we will work strictly in $\mathcal{L}^{2}(\Omega,\mathcal{F},P)$.
Key facts about Hilbert spaces
We state in the following a few useful results, proofs can be found in any standard text on functional analysis.
Theorem (Canonical norm).
Any inner product space $\mathcal{X}$ can be endowed with the norm
$\|x\|:={\sqrt{\langle x,x\rangle}},\quad x\in\mathcal{X}.$
This is called the inner product norm, or canonical norm, or the norm induced by the inner product
Definition.
[…]
Definition.
$x,y\in\mathcal{X}$ are called orthogonal if $\langle x,y\rangle=0$
Theorem.
For every point $x$ in a Hilbert space $\mathcal{H}$ and every nonempty closed convex subset $C\subseteq H$ there exists a unique vector $y\in C$ for which $\lVert x-y\rVert$ is equal to $\delta:=\underset{c\in C}{\inf}\|x-c\|$
Theorem (Hilbert projection theorem).
Let $\mathcal{H}_{1}\subset\mathcal{H}$ be a subspace of a Hilbert space $\mathcal{H}$. Then, $\forall x\in\mathcal{H},\ \exists$ a unique vector $\mathcal{P}x\in\mathcal{H}_{1}$ such that $\forall z\in\mathcal{H}_{1},\ \langle x-\mathcal{P}x,z\rangle=0$. $\mathcal{P}x$ is called the projection of $x$ onto $\mathcal{H}_{1}$. In addition, $\mathcal{P}x$ is the element of minimum distance from $\mathcal{H}_{1}$
Proposition.
The operator $\mathcal{P}$ above has the following properties:
  • $\mathcal{P}$ is a linear contraction
  • $\mathcal{P}$ is idempotent
[…]
Proposition.
If $C$ is a closed vector subspace of a Hilbert space, then it is a Hilbert subspace. In addition, every finite dimensional vector subspace of an Hilbert space is closed
3.2 The nature of mathematical models: definitions
We have now the motivation and have available all the necessary structures for a formal treatment of what mathematical models for the applied sciences are. We refer to the above definition of $\mathcal{H}$.
Definition.
Given a random variable of interest or objective random variable, $\boldsymbol{X}^{o}$, possibly supposed to be related to some random variable $\boldsymbol{U}$, a model of $\boldsymbol{X}^{o}$ is an operator $\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{U}):\mathcal{H}\times\mathbb{R}^{q}\rightarrow\mathcal{H}$, whose arguments are some parameter $\boldsymbol{\theta}\in\mathbb{R}^{q}$ together with the predictor random variable(s) $\boldsymbol{U}$ (if any), and whose value is another random variable $\boldsymbol{X}=\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{U})$.
Remark.
That $\boldsymbol{X}$ is a random variable follows from the Doob-Dynkin lemma. If $\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{U})$ is to be a good model, then $\boldsymbol{X}$ should be close to $\boldsymbol{X}^{o}$ in the metric of $\mathcal{H}$.
Definition.
Having fixed the predictor(s) $\boldsymbol{U}$ , we will typically consider the relationship between the parameter and the value of the operator, the model map $\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\theta})$  .
Remark.
Notice that the above is a slight abuse of notation: since we are typically interested in the relationship between the parameter $\boldsymbol{\theta}$ and the value $\boldsymbol{X}$ of the model $\boldsymbol{x}(\boldsymbol{\theta},\boldsymbol{U})$, having for all intents and purposes fixed $\boldsymbol{U}$, we just consider $\boldsymbol{X}$ as a function of $\boldsymbol{\theta}$. To keep the two concepts distinct however, we use the two different names "model" and "model map".
Definition.
Having fixed $\boldsymbol{U}$ and given sufficient regularity of the model, the set of values =$\mathcal{M}=\{\boldsymbol{X}(\boldsymbol{\theta}),\boldsymbol{\theta}\in\boldsymbol{\Theta}\}$, indexed by the parameter, is the model manifold, $\mathcal{M}\subset\mathcal{H}$
3.3 Linear and nonlinear models
If the hypothesized model is linear, e.g. $\boldsymbol{X}(\boldsymbol{\theta})=\boldsymbol{U}\,\boldsymbol{\theta}$ , then it is easy to see that $\mathcal{M}$ is a vector space with elements in $\mathcal{H}$, hence a subspace of the Hilbert space $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathcal{F},\mathbb{P})$.
As a subspace of a Hilbert space, $\mathcal{M}$ in this case inherits its Hilbert space structure (norm, inner product), and a projection operator is guaranteed. Consequently, the point of projection $\hat{\boldsymbol{X}}=\mathcal{P}_{{\mathcal{M}}}(\boldsymbol{X}^{o})$ is the one of minimal distance from $\boldsymbol{X}^{o}$ (in the norm of $\mathcal{H}$), is unique and    ($\boldsymbol{X}^{o}-\hat{\boldsymbol{X}}$)   is orthogonal to all elements of $\mathcal{M}$.
If on the other hand the model is nonlinear, e.g. $\boldsymbol{X}(\theta)=\sin(\theta\boldsymbol{U})$  , then $\mathcal{M}$ is not in general convex, much less a subspace of $\mathcal{H}$. In our example in fact :=$\mathcal{M}:=\{\boldsymbol{X}(\theta)\}$ is not convex since for $\theta^{1},\theta^{2}\in\boldsymbol{\Theta},\theta^{1}\neq\theta^{2}$ we have $X(\theta^{1})+X(\theta^{2})\notin\mathcal{M}$. Hence $\mathcal{M}$ is not a vector space, not a subspace of the Hilbert space $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathcal{F},\mathbb{P})$, no projection operator $\mathcal{P}_{{\mathcal{M}}}$ can be defined and no unique “best" estimate $\hat{\boldsymbol{X}}$ can be obtained.
However, assume $\exists$ open $\,\mathcal{U}\subset\mathcal{M}$ and $\hat{\boldsymbol{X}}=\boldsymbol{X}(\hat{\boldsymbol{\theta}})\in\mathcal{U}$ such that $\forall\boldsymbol{X}\in\mathcal{U}$, $\boldsymbol{X}\neq\hat{\boldsymbol{X}}$ it is $\mathrm{d}(X^{o},\hat{X})<\mathrm{d}(X^{o},X)$. Hence, we assume we have an isolated point $\hat{X}\in\mathcal{M}$ locally closest to $X^{o}$. Under this hypothesis, that a random variable $\hat{\boldsymbol{X}}$ exists, of minimal distance from $\boldsymbol{X}^{o}$ within a neighborhood (in the topology on $\mathcal{H}$ induced by the norm), then we could consider working on a local linear approximation to the manifold $\mathcal{M}$ at $\hat{\boldsymbol{X}}$, that is we could consider constructing the tangent space to the manifold $\mathcal{M}$ at that locally optimal point, $T_{\hat{\boldsymbol{X}}}\mathcal{M}$. To do so, we need to use the concept of the tangent space to a generic topological manifold at one of its points. Any good introductory text in differential geometry (e.g. [17, 18, 19]) will provide the relevant definitions and proofs. Notice, again, that in general $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ is not a subspace of $\mathcal{H}$, but rather an affine space. However, the Hilbert projection theorem extends naturally to affine subspaces by translation, see Appendix 5.1 for details: here we will directly deal with the local approximation $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ to $\mathcal{M}$ at $\hat{\boldsymbol{X}}$, calling, with a slight abuse of language, the tangent space $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ to $\mathcal{M}$ (which is a bona-fide vector space in its own right) a “subspace” of $\mathcal{H}$, see Fig.LABEL:fig:NonLinearHilbertReparametrized.
We can build $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ by considering the set of velocities, i.e. of differential operators, which map functions in $C^{\infty}(\mathcal{M})$ , $f:\mathcal{M}\rightarrow\mathbb{R}$ to their directional derivatives in the direction of curves $\gamma:\mathbb{R}\rightarrow\mathcal{M}$, $T_{\hat{\boldsymbol{X}}}\mathcal{M}=\{\vartheta_{\gamma,\hat{\boldsymbol{X}}}$ such that $\,\hat{\boldsymbol{X}}=\gamma(\lambda)\in\mathcal{U}\subset\mathcal{M}$ and $\forall f\in C^{\infty}(\mathcal{M})$ it is $\,\vartheta_{\gamma,\hat{\boldsymbol{X}}}=(f\circ\gamma)^{\prime}|_{\lambda}\}$. We can in general define a chart-induced basis of the tangent space: in the present case we will compute the basis for $T_{\hat{\boldsymbol{X}}}\mathcal{M}$ induced by the canonical chart $\boldsymbol{X}^{-1}$:
$\mathcal{U}$$\mathbb{R}^{q}$$\boldsymbol{X}^{-1}$$\boldsymbol{X}$
Clearly we must require sufficient regularity in the model map $\boldsymbol{X}$ to do this, such as being continuously differentiable with a continuously differentiable inverse. Notice that $\boldsymbol{X}^{-1}$ is a chart, indeed a global chart, mapping the whole of $\mathcal{M}$ to $\mathbb{R}^{q}$, since $\mathcal{M}$ by definition is generated by $\boldsymbol{X}(\boldsymbol{\Theta})$, with $\boldsymbol{\Theta}\in\mathbb{R}^{q}$. We can also say, therefore that:
Lemma.
The inverse of a model map is a chart map.
It is now necessary to clarify the correspondence between the (differential) operators in $T_{\hat{\boldsymbol{X}}}(\mathcal{M})$ and the random variables that ought to belong to the (affine) subspace of $\mathcal{H}$ it defines.
We can do this by using the “canonical" chart from the manifold to $\mathbb{R}^{q}$ induced by the parametrization of the model. As is detailed in Appendix LABEL:sec:AffineTangentToM We can express the directional derivative of a function $f$ in the direction of a curve $\gamma$ at a point $p=\gamma(\lambda)$, with respect to a specific chart map $\mu$, as the dot product of the components of the derivative of the curve representation under the chart $\gamma_{\mu}$ at $\lambda$, times the application of the operator $\frac{\partial}{\partial\mu}|_{p}=(\cdot\circ\mu^{-1})|_{p_{\mu}}$ to the function $f$.
For our specific application we can use as basis of the tangent space the basis induced by the chart map $(\boldsymbol{X}^{-1})$:
$\{(\cdot\circ((\boldsymbol{X}^{-1})^{-1})^{i})^{\prime}|_{\hat{\boldsymbol{X}}}\}_{i=1,...,q}=\{(\cdot\circ\boldsymbol{X}^{i})^{\prime}|_{\hat{\boldsymbol{X}}}\}_{i=1,...,q}\,\,.$
Denoting $\dot{\boldsymbol{X}}:=\boldsymbol{X}^{\prime}$ we have that $(\boldsymbol{X}^{i})^{\prime}=(\dot{\boldsymbol{X}})^{i}=\dot{\boldsymbol{X}}^{i}$     hence the tangent space is the span $T_{\hat{\boldsymbol{X}}}{\mathcal{M}}=\mathcal{S}\{\dot{\boldsymbol{X}}^{i}\}_{i=1,...,q}$    WRT the canonical model chart map $(\boldsymbol{X}^{-1})$. This means that for any $\eta\in\mathbb{R}^{q}$, $(\eta\cdot\dot{\boldsymbol{X}}+\hat{\boldsymbol{X}})\in T_{\hat{\boldsymbol{X}}}\mathcal{M}\subset\mathcal{H}$ and $(\eta\cdot\dot{\boldsymbol{X}}+\hat{\boldsymbol{X}})$ is a random variable in the (affine) tangent space to the model manifold $\mathcal{M}$ at $\hat{\boldsymbol{X}}$.
An affine subspace $T_{\hat{\boldsymbol{X}}}\mathcal{M}\subset\mathcal{H}$ of random variables has thus been constructed using the model canonical chart map $\boldsymbol{X}^{-1}$. The Hilbert projection theorem applies and the random variable $\boldsymbol{X}^{o}$ has a unique projection onto $\hat{\boldsymbol{X}}\in T_{\hat{\boldsymbol{X}}}\mathcal{M}$ , of global minimal distance from $T_{\hat{\boldsymbol{X}}}\mathcal{M}$, such that $\forall\,\boldsymbol{X}\in T_{\hat{\boldsymbol{X}}}\mathcal{M}$ we have $(\boldsymbol{X}^{o}-\hat{\boldsymbol{X}})\,\perp(\boldsymbol{X}-\hat{\boldsymbol{X}})$.
3.4 Relationship with conditional estimation
As already mentioned in the Introduction, much work has been done on the Hilbert space of all finite-variance random variables on a probability space $(\Omega,\mathcal{F},P)$, i.e. on all $\mathcal{F}$-measurable functions on $\Omega$. Estimation of a random variable $\boldsymbol{X}^{o}\in\mathcal{H}$ conditional on a sub-$\sigma$-algebra $\mathcal{F}_{\mathcal{G}}\subset\mathcal{F}$, i.e. assuming a lesser amount of information than that provided by the full $\sigma$-algebra $\mathcal{F}$, is shown to consist in the projection of the original random variable $\boldsymbol{X}^{o}$ onto the subspace $\mathcal{G}\subset\mathcal{H}$ of $\mathcal{F_{G}}$-measurable random variables. In this case we do have a (proper) subspace, the Hilbert projection theorem holds and the projection of $\boldsymbol{X}^{o}\in\mathcal{H}$ onto $\mathcal{G}$ is the conditional estimate $\hat{\boldsymbol{X}}$ (conditional on $\mathcal{F_{G}}$).
Notice however that the situation described in the present work is different: even in the case of linearity, the subspace of interest is generated by a functional form of the model, linear with respect to the parameter values, not by the restriction of information on the outcome characterized by an appropriate sub-$\sigma$-algebra. The difference between the two situations is obvious when considering a nonlinear functional form of the model.
3.5 Relationship with Information Geometry
The well-developed field of Information Geometry deals with spaces of distributions, wherein certain (parametrized) families are embedded manifolds. Interesting properties can be derived for some such families (e.g. the exponential family generates a "flat" manifold).
Very interesting facts emerge from the study of such spaces of distributions. For example, denote by $\mathcal{P}$ the space of probability distributions, and denote by $\mathcal{F}\subset\mathcal{P}$ the manifold representing a family of probability distributions indexed by some parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{q}$, $F(\boldsymbol{\theta})\in\mathcal{F}$, which family we wish to use as a possible “model” for our data-generating process. We assume that the data are actually generated by an unknown distribution $P\in\mathcal{P}$, with, in general $P\notin\mathcal{F}$. If we observe data $\boldsymbol{x}=\{x_{1},...,x_{n}\}$ we can build an empirical distribution $\tilde{P}$ out of them. We could then hypothesize that the “best” value of $\boldsymbol{\theta}$ is the value $\hat{\boldsymbol{\theta}}$ that minimizes some kind of distance or divergence from $\tilde{P}$ to $F(\boldsymbol{\theta})$: in the case the Kullback-Leibler divergence $D_{KL}$ of $F(\boldsymbol{\theta})$ from $\tilde{P}$ is considered, we obtain the Maximum Likelihood Estimator $\hat{\boldsymbol{\theta}}_{ML}$.
Now, there is a (biunivocal $P$-almost-everywhere) relationship between random variables on $(\Omega,\mathcal{F},P)$ and their distributions, in the sense that each random variable induces a distribution and conversely we can define a random variable from a distribution, two random variables with the same distribution being equal except on $P$-null sets. In other words, when we model a physical phenomenon we are working in $\mathcal{H}$, and when we work on the properties of estimators we use results established in $\mathcal{P}$, but there is an almost-everywhere 1-1 correspondence between the two.
3.6 Mapping $\mathcal{H}$ to $\mathbb{R}^{n}$
We need now to formalize the intuitively obvious relationship between the theoretical construction of the abstract model in the Hilbert space of finite-variance random variables $\mathcal{H}$ and the actual computational, numerical analysis of the results carried out with statistical techniques in case space $\mathbb{R}^{n}$. A random variable is defined as a measurable function from $(\Omega,\mathcal{F},P)$ to $\mathbb{R}^{n}$, say
$Y\in\mathcal{H}\quad\Rightarrow\quad Y:\Omega\rightarrow\mathbb{R}^{n}\,\,,\,\,\,Y:\omega\mapsto r\,.$
However, we might equally well define an event $\omega\in\Omega$ as a map from $\mathcal{H}$ to $\mathbb{R}^{n}$:
$\omega\in\Omega\quad\Rightarrow\quad\omega:\mathcal{H}\rightarrow\mathbb{R}^{n}\,\,,\,\,\,\omega:Y\mapsto r\,.$
In other words, we may define
$\omega(Y):=Y(\omega)=r\in\mathbb{R}^{n}$ (4)
Experimental sampling is the concrete act of picking an event $\omega$ from the sample space $\Omega$. This act determines the specific realization $\boldsymbol{x}^{o}$ of the random variable $\boldsymbol{X}^{o}$. As a consequence, this act also determines the realization $\hat{\boldsymbol{x}}$ of $\hat{\boldsymbol{X}}$, and so on. All random variables considered in the theoretical construction of the model are thus naturally mapped by $\omega$ to their respective numerical realizations. We see then that the event $\omega$ can be interpreted as the necessary map from $\mathcal{H}$ to $\mathbb{R}^{n}$, see Fig.4.
Figure 4: Sampling as a map from $\mathcal{H}$ to $\mathbb{R}^{n}$.
4 Discussion and Conclusions
Recent literature addresses similar problems (in particular the characterization of the projection) in the case of conditional expectation [4] and in the case of information geometry [6, 7, 10, 8, 9]. Specifically, in information geometry a set of probability distributions indexed by a finite number of parameters is shown to be a (smooth) manifold and is equipped with a Riemannian structure and with conjugate connections, which make the manifold ‘flat’ and generalize the concept of Euclidean Space [8, 6]. This ‘flat’ structure can be obtained in various ways, some of which are discussed in the papers of Amari et al. and Nielsen [9, 8, 6, 11]. One way of doing this is through Bregman divergence, which induces a structure on a manifold $\mathcal{M}$ that generalizes the Pythagorean Theorem and consequently guarantees the existence and uniqueness of the projection of a point $P$ onto the manifold $\mathcal{M}$ [9, 8]. A similar approach to Amari (also formulating a projection theorem) is proposed by Brigo [10] using a direct $\mathcal{L}^{2}$ metric. We note that, as stated above, the theory of conditional expectation as a projection onto an appropriate subspace of the Hilbert space $\mathcal{H}$ of random variables is not related to the current discussion because the underlying manifold is different. As regards information geometry, there is a clear connection since any random variable induces a probability distribution, but the identification of the underlying manifold as determined by the mathematical model of the (biological) experiment, as well as the identification of the map between the space $\mathcal{H}$, where the modeling is abstractly done, and $\mathbb{R}^{n}$, where the computations are carried out, to the best of our knowledge have not yet been formalized.
In conclusion, a mathematical model is an operator yielding a random variable $\boldsymbol{X}$ approximating the objective random variable $\boldsymbol{X}^{o}$ in $\mathcal{H}$. The event $\omega$ maps $\mathcal{H}$ onto case space $\mathbb{R}^{n}$. Estimation for linear models in case space consists in projecting the observed value $\boldsymbol{x}^{o}$ onto the prediction surface $\mathcal{S}$, and corresponds via $\omega$ to projecting $\boldsymbol{X}^{o}$ onto the model manifold $\mathcal{M}$. For nonlinear models a local optimum $\hat{\boldsymbol{x}}\in\mathcal{S}$ can be found numerically, and is the image under $\omega$ of a locally optimal random variable $\hat{\boldsymbol{X}}\in\mathcal{M}$. Asymptotic confidence regions can be computed on $\mathcal{S}$ for linear models. For nonlinear models approximate asymptotic confidence regions can be computed on the tangent space $T_{\hat{\boldsymbol{x}}}(\mathcal{S})$ (constructed via computation of the Jacobian at the local optimum), which corresponds to the theoretical tangent space of differential operators $T_{\hat{X}}(\mathcal{M})\subset\mathcal{H}$.
5 Appendices
5.1 Extension of Hilbert’s theorem to affine spaces
The classical Hilbert projection theorem states that if $\mathcal{H}$ is a Hilbert space and $\mathcal{V}\subset\mathcal{H}$ is a nonempty closed convex subset, then for every $x\in\mathcal{H}$, there exists a unique point $p\in\mathcal{V}$ such that
$\|x-p\|=\inf_{y\in\mathcal{V}}\|x-y\|.$ (5)
In particular, when $\mathcal{V}$ is a closed linear subspace, the projection $p$ satisfies
$x-p\perp\mathcal{V}.$
Now, let $\mathcal{H}$ be a Hilbert space, and let $\mathcal{S}$ be an affine subspace, meaning there exists a closed linear subspace $\mathcal{V}$ and a fixed point $a\in\mathcal{H}$ such that
$\mathcal{S}=a+\mathcal{V}=\{a+v:v\in\mathcal{V}\}.$ (6)
Theorem.
Let $\mathcal{H}$ be a Hilbert space, $\mathcal{V}\subset\mathcal{H}$    a nonempty closed subspace and $\mathcal{S}=a+\mathcal{V}$ a nonempty closed affine subspace. Then for every $x\in\mathcal{H}$, there exists a unique $p\in\mathcal{S}$ such that
$\|x-p\|=\inf_{y\in\mathcal{S}}\|x-y\|.$
Moreover,
$x-p\perp\mathcal{V}.$
Proof.
Since $\mathcal{S}$ is an affine subspace, we can write $\mathcal{S}=a+\mathcal{V}$. Define the translation $x^{\prime}=x-a$, so that our problem reduces to projecting $x^{\prime}$ onto the linear subspace $\mathcal{V}$. By the Hilbert projection theorem for linear subspaces, there exists a unique $v_{0}\in\mathcal{V}$ such that
$\|x^{\prime}-v_{0}\|=\inf_{v\in\mathcal{V}}\|x^{\prime}-v\|.$ (7)
This implies that $x^{\prime}-v_{0}$ is orthogonal to $\mathcal{V}$, i.e.,
$x^{\prime}-v_{0}\perp\mathcal{V}.$ (8)
Setting $p=a+v_{0}\in\mathcal{S}$, we obtain
$x-p=(x^{\prime}-v_{0})\perp\mathcal{V}.$ (9)
Thus, $p$ is the unique closest point in $\mathcal{S}$ to $x$, and the orthogonality condition still holds relative to the associated direction subspace $\mathcal{V}$. ∎
Acknowledgments
The work of Prof. Andrea De Gaetano was supported by the Distinguished Professor Excellence Program of Óbuda University, Budapest Hungary.